OpenStudy (anonymous):

a radio active substance has a half-life of 200 years. We currently have 15 grams of the substance. How much of the substance will still be radioactive 80 years from now?

OpenStudy (amistre64):

-ln(2)/200 = rate of decay if i see it correctly

OpenStudy (amistre64):

A = 15 e^(-ln(2)/200 * 80) perhaps?

OpenStudy (anonymous):

How long will it take until 1/2 grams of the substance remains radioactive?

OpenStudy (amistre64):

.5 = 15 e^(-kt) .5/15 = e^(-kt) ln(.5/15) = -kt t = ln(.5/15)/(-ln(2)/200) maybe?

OpenStudy (amistre64):

981 years or so .... i think

OpenStudy (anonymous):

this should help me thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

a one time deposit is made into an account that pays an apr of 8.5% compounded continuously. How long does it take for the worth of the account to triple?